OpenStudy (anonymous):

can someone help me find the solution to: arctan(2sin(pi/3)) a walkthrough would help!

OpenStudy (anonymous):

find \(\sin(\frac{\pi}{3})\) then double it then take the arctangent of the result

OpenStudy (anonymous):

yeah i did that but the answer I'm getting isn't right apparently

OpenStudy (anonymous):

oh wait nvm i found what i was doing wrong